### v1

# Adversarial Agent Collaboration for C to Rust Translation

Abstract
Translating C to memory-safe languages, like Rust, prevents critical memory safety vulnerabilities that are prevalent in legacy C software. Existing approaches for C to safe Rust translation, including LLM-assisted ones, do not generalize on larger (>500>500 LoC) C codebases because they depend on complex program analyses that frequently break. In this work, we present ACToR (Adversarial C To Rust translator), a simple LLM agent-based approach. Inspired by GANs, ACToR pits a generator agent against a discriminator agent, which collaborate to iteratively generate a Rust translation. On each iteration, the translator agent synthesizes and refines a Rust translation to pass an existing suite of tests, and then the discriminator agent finds new failing tests. We demonstrate that ACToR translates all of the 63 real-world command line utilities considered in our benchmarks, which have an average size of 485 lines of code, and it achieves over 90% test pass rate with zero human intervention. To our knowledge, it is the first such system that reliably translates C programs of this scale. Furthermore, ACToR improves translation correctness by up to 18.9% compared to baseline, non-adversarial approaches.
1 Introduction
Memory safety vulnerabilities in C/C++ code constitute a large fraction of security vulnerabilities each year. Microsoft reported 70% of their CVEs are due to memory safety issues (Microsoft, 2019a; b). In response, there has been an increasing demand to translate memory-unsafe legacy code to modern memory-safe languages like Rust from industry and governments (DARPA, 2025), thereby ensuring full memory safety by design. Unfortunately, given the scale of legacy code that permeates our software systems today, manual translation of (millions of lines of) existing C code is a formidable and practically infeasible task. A reasonable translation from C to Rust must use safe and idiomatic Rust programming abstractions, which are checked mostly at compile time and avoid runtime performance overhead, unlike memory safety checking directly in C (Nagarakatte et al., 2010; 2009; Orthen et al., 2024; Serebryany et al., 2012).
Existing approaches for automated translation come in two flavors: rule-based and LLM-assisted. Rule-based translators, like c2rust (Immunant, ) and its derivatives (Zhang et al., 2023; Emre et al., 2023; Ling et al., 2022), produce unidiomatic and unsafe Rust, which limits their usefulness in practice. LLM-assisted translators produce more idiomatic and safe Rust; however, current works along this line fail to generalize to large (>500>500 lines of code) C projects. No prior work on LLM-assisted translation has demonstrated their approach successfully translating more than one large C project, and several reports require manual intervention owing to this brittleness (Shetty et al., 2024; Cai et al., 2025). This brittleness is largely due to their dependency on complex program analyses that break on unseen C programs.
In this work, we aim to advance the state-of-the-art in automated translation and turn our attention towards LLM agents. LLM agents have shown promising capabilities in solving generic but complex software engineering tasks (Jimenez et al., 2023) owing to their ability to iteratively interact with development environments, utilizing tools like code editors, compilers, and unit testing to detect and correct errors to improve results (Chen et al., 2023; Dong et al., 2024; Huang et al., 2023; Qian et al., 2024). In the context of automated translation, a straightforward agentic setup is as follows —- the agent has access to the development environment including the Rust compiler, and proceeds by translating the C code to Rust, and iteratively refines it by incorporating feedback from the compiler (in case the translated Rust program does not compile) or from the test inputs on which the translated code provides an output that differs from the output of the original C program. Test suites serve as a useful and pragmatic proxy for the more rigorous, but largely elusive and impractical alternative of full-fledged formal verification for semantic equivalence between the cross-language code, and have unanimously been deployed in prior works (Yang et al., 2024; Eniser et al., 2024; Wang et al., 2025; Zhang et al., 2025).
While the ability to access development environment without a priori fixed workflow is a promising first step towards more reliable and realistic translations, we find, rather unsurprisingly, that simplistic code agents like the one described above nevertheless produce incorrect translations; the behavior of the translated Rust code they produce diverges significantly from that of the original C code on inputs absent from the test suite. As a result, such a vanilla agentic setup requires frequent developer intervention and offers little to no improvement over traditional LLM-based translation techniques. At a high level, a simplistic agentic setup, like the one we just described, lacks the key ability of generalization — it produces a translation that is overfitted to the (typically small) test suite available to it, but often cannot ensure expected behavior on inputs absent from the suite. Unfortunately, as our evaluation shows, simple fixes such as by increasing the test coverage on the C source, which we achieve by augmenting agents with standard test coverage tools (GNU, 2025), continue to remain ineffective in enhancing the quality of translations.
In this work, we introduce a new agentic setup for program translation designed to overcome the limited generalization observed in simpler agent frameworks. Inspired by the generator–discriminator paradigm popularized by GANs (Goodfellow et al., 2020), we employ two collaborating agents: a translator, which proposes candidate translations from C to Rust, and a discriminator, which actively searches for evidence( in the form of counterexamples) that the translation diverges from the original. Whenever the discriminator identifies such adversarial inputs, these are fed back to the translator, which uses them to refine subsequent translations. Through this iterative process, the translator learns to produce Rust programs that not only pass the original test cases but also withstand adversarial scrutiny, leading to translations that are more robust and semantically faithful to the original C program. We refer to this adversarial translator–discriminator framework as ACToR, for Adversarial C To Rust, as demonstrated in Figure 1.
We demonstrate the effectiveness of ACToR on two sets of benchmarks: A micro-benchmark set of $6$ C programs considered in recent work, and, a much larger macro-benchmark set of $57$ C programs from the standard BSDCoreUtils. These $63$ real-world C programs range from $7$ to $5,469$ lines of code (LoC), with a median of 485, and totaling over 30,000. ACToR automatically produces Rust translations for all benchmarks that pass more than 90% of tests on average. To our knowledge, it is the first system to yield translations with zero human intervention and with high pass rates for this scale of programs. In addition, we show that ACToR’s adversarial design choice improves pass rates by up to 18.9% compared with the baseline, and it also improves the quality of tests generated, as measured by their ability to find discriminating behavior.
Figure 1: High-level overview of ACToR. The Translator and the Discriminator agents update the translation and the tests in turn to iteratively improve the correctness of the translated program.
2 Related Work
Automated translation of C to Rust is a relatively recent approach towards achieving full memory safety (Emre et al., 2021). The desired goal of such translation is to generate safe [^1] The use of unsafe keyword in Rust allows mixing unchecked and possibly memory unsafe code in otherwise safe Rust Rust code corresponding C code, i.e., Rust code that uses only idiomatic Rust primitives that the compiler can statically verify or trust to be memory safe. While several prior works have tried, none have worked reliably even for programs of moderate size (e.g., about $500$ LoC) for varied reasons (Li et al., 2024).
2.1 Program Translation (C to Rust)
Existing works on C-to-Rust translation can primarily be categorized as rule-based or LLM-based techniques. Rule-based techniques such as c2rust (Immunant, ) and corrode (Sharp, ) opt for mechanical, line-by-line, rewriting of the C code, often resorting to raw pointers and unsafe Rust. Later efforts explored ways to infer Rust types from C code, either by compiler feedback (Emre et al., 2023; Hong & Ryu, 2025b), SMT solving (Zhang et al., 2023), data-flow graphs (Xu & Huang, 2025), or syntactic rewriting (Ling et al., 2022), and specialized analyses targeted idioms like tagged unions (Hong & Ryu, 2024b), output parameters (Hong & Ryu, 2024a), and file APIs (Hong & Ryu, 2025a). Beyond general C, Low* compilation to Rust (Fromherz & Protzenko, 2024) demonstrates scalable translation in restricted settings, though such techniques remain infeasible for general-purpose C due to expressiveness and unmodeled APIs.
LLM-based approaches use an LLM to propose a translation (Lachaux et al., 2020). Recent work augments models with different kinds of tools. These include test generation tools (Eniser et al., 2024; Yang et al., 2024), dynamic pointer analyses (Shetty et al., 2024), or multi-modal prompting (Nitin et al., 2024). Function-level decomposition has been explored to improve scalability (Shiraishi & Shinagawa, 2024; Cai et al., 2025; Ou et al., 2025), while agentic workflows combine LLMs with static analysis (Zhou et al., 2025; Farrukh et al., 2025). Neural methods yield more idiomatic code, yet often fall short of preserving exact input/output behavior. Moreover, current work fails to generalize on larger C programs. Studies echo these concerns: decomposition introduces inconsistencies (Li et al., 2024), and LLMs remain prone to subtle errors (pan, 2024).
Work beyond C-to-Rust highlights related strategies. Go-to-Rust translation explored scalable decomposition (Zhang et al., 2025); TransMap (Wang et al., 2023b) and DuoGlot (Wang et al., 2023a) used semantic checking or guided search; skeleton-based methods modeled programs as composable fragments (Wang et al., 2025). Classical efforts translated across other pairs (C++→Java (Buddrus & Schödel, 1998; Malabarba et al., 1999; Beevi et al., 2014), Java→Python (Lachaux et al., 2020; Ibrahimzada et al., 2025)). Meanwhile, verified lifting has advanced in domain-specific contexts, including numerical kernels (Kamil et al., 2016), tensor operators (Zhan et al., 2024; Qiu et al., 2024), and grammar-driven rewriting (Cordy, 2006), with early LLM-based attempts (Bhatia et al., 2024). These exemplify the correctness ideal but are not yet generalizable to C.
Overall, there is growing consensus in favor of techniques that combine rule-based precision with neural readability. Though, to the best of our knowledge, none of the existing techniques is automatic—they either primarily work on small pieces of code or have a complex implementation, often requiring human interventions when aiming for functioning translations of whole C programs. In contrast, a prime focus of our work is the design of a simple framework that leverages LLM agents for C to Rust translation which is easier to implement but also robust enough to automatically translate whole programs to functioning safe Rust code.
2.2 Collaborative LLMs & Agents
Many works propose multi-LLM and multi-agent systems, where the LLMs or agents collaborate to produce a final answer. Many works on code generation propose specific roles, such as coder, tester, or analyst, which collaborate to write code (McAleese et al., 2024; Chen et al., 2023; Dong et al., 2024; Huang et al., 2023; Qian et al., 2024; Islam et al., 2024). Collaborative approaches have been used outside of code generation as well. Many works propose self-reflection techniques (Shinn et al., 2023; Madaan et al., 2023; Dhuliawala et al., 2023), sometimes aided by tools (pan, 2024; Gou et al., 2024; Dong et al., 2025), where the LLM first proposes an answer, then reflects on its answer to improve accuracy. Similar works propose multi-agent debate architectures (Chan et al., 2024; Du et al., 2023; Liang et al., 2023; Hong et al., 2024; Wu et al., 2024; Liu et al., 2024) where multiple answers are proposed, then the LLM(s) debate the correct answer. Several works propose frameworks for agent collaboration (Hong et al., 2024; Wu et al., 2024; Liu et al., 2024).
The key difference between ours and prior work is in the high-level principle: To reliably improve the translated code, we leverage the C program as the oracle in the construction of a discriminator, which automatically identifies correctness gaps and provides iterative feedback to the translator.
3 ACToR Design
We consider a realistic setup where C project is required to be translated to safe Rust. We explain how our adversarial translation approach solves this task with no human intervention beyond providing the original C program and a set of initial seed test cases.
3.1 Code Translation Task
Given a source C program $c$, the goal of the translation is to find a functionally correct Rust code $rs$ that has equivalent external behaviors as the source C program on all valid inputs. External behaviors include standard output, error messages, file modifications, and so on. More specifically, we want the translated code to satisfy the boolean requirement $\texttt{IsEq}(c,rs)$, where IsEq is an oracle that takes a C program and a Rust program as input, and determines whether they have equivalent behaviors on the valid universe $U$ of inputs that do not violate memory safety. This objective naturally captures the desired requirement that the translated Rust program has the same behavior as the C program on safe inputs. On the other hand, memory-safe handling of malicious inputs is guaranteed by the Rust compiler itself.
A naive approach to obtaining a Rust translation is to prompt the LLM agent with the C code $c$ and a small finite set $T\subset U$ of inputs sampled from the safe input universe $U$ with the goal of computing a Rust program that matches the behaviour of $c$ on all inputs from the sampled set $T$. Unfortunately, such a naive method does not produce a satisfactory translation. The space of valid inputs is typically infinite for most programs. Indeed, we empirically observed that with only a limited number of input instances tested one-shot by the agents, the produced translation does not generalize to unseen test cases beyond the limited set $T$.
We remark that a perfect oracle IsEq does not exist in practice. There is no automated tool that practically works in determining whether the given C code and Rust are equivalent in the input universe, or provides counterexamples when they mismatch. On the contrary, what we have is ‘IsEq*’, which takes as input both the C and the Rust code, together with a concrete input $t$, tells whether they mismatch on this input $t$. This intuitively means we run the C code and Rust code on the same input, and check if they result in the same effect. The problem now becomes how to satisfy IsEq with only IsEq* available.
The next most natural solution is to continually sample new test cases from the input universe $U$ and use them to improve the translated code iteratively. However, this approach is not particularly efficient and does not help much in terms of improving the equivalence on $U$. As the number of tests taken increases, the probability of sampling an input that re-explores a previously tested behavior increases. The incremental improvement per new test becomes modest as we proceed.
3.2 Adversarial Translation
We propose ACToR, which leverages the adversarial collaboration between two agents. The idea behind ACToR comes from GAN, where two opponents are included to estimate a data distribution implicitly. ACToR consists of two collaborative agents: the translator agent and the discriminator agent. The discriminator aims to surface the mismatches in translated code, while the translator aims to improve the translation to evade detection by the discriminator. The two agents work in turn, iteratively improving the translation. The detailed algorithm is shown in the Algorithm 1.
Algorithm 1 ACToR: Adversarial C to Rust Translation
1:C program $c$, seed test set $T_{0}$, test input universe $U$ (implicit)
2:Params: MaxIter (of outer loop), TestBatchSize, RetryLimit (of inner loops)
3:Final translation $rs^{*}$ and final test set $T^{*}$
4:while Not PassesSuite($c,rs,T_{0}$) do $\triangleright$ initial translation on seed tests
5:  $rs\leftarrow$ Translator($c,\varnothing,T_{0}$)
6:end while
[…]
8:for $k=1$ to MaxIter do
9:  $numRetry\leftarrow 0$
10:  repeat
11:   $Batch\leftarrow$ Discriminator($c,rs,n=\text{TestBatchSize}$)
12:   $IsValid\leftarrow\big(\forall t\in Batch:\;\textsc{ValidTestFormat}(t)\land\textsc{SanityOk}(t,c)\big)$
13:   $IsDisc\leftarrow\big(\exists t\in Batch:\;\neg\texttt{IsEq}^{*}(c,rs,t)\big)$
14:   $numRetry\leftarrow numRetry+1$
15:   $IsReachingLimit\leftarrow(numRetry\geq\text{RetryLimit})$
16:  until $(|Batch|=\text{TestBatchSize}~\land~IsValid~\land~IsDisc)~\lor~IsReachingLimit$
17:  $T\leftarrow T\cup Batch$
18:  $numRetry\leftarrow 0$
19:  while (Not PassesSuite($c,rs,T$)) $\lor~numRetry\leq\text{RetryLimit}$ do
20:   $numRetry\leftarrow numRetry+1$
21:   $rs\leftarrow$ Translator($c,rs,T$)
22:  end while
23:end for
24:return rsrs as $rs^{*}$, $T$ as $T^{*}$
ACToR takes as input the C code $c$, the universe $U$, the seed test set $T_{0}$, together with some iteration parameters such as the number of times the translator and discriminator interact (MaxIter) before ACToR outputs the final translated program. ACToR then conducts iterative translation on the source C code and outputs both the Rust translation rsrs* and the final test set TT*. During iterative improvement, ACToR maintains an append-only test set $T$. In the beginning, the translator agent generates the initial translation and ensures it passes all the seed tests (lines 1-3) before the adversarial iterative loop (involving the discriminator) begins. Next, in each iteration that follows, the discriminator agent reads the C code $c$ and the current Rust code $rs$, and generates test cases on which C and Rust outputs disagree (lines 7-13). The discriminator is structured so that (1) all the test cases abide by the correct format and also pass basic sanity checks when running them on the source C program (line 9); and that (2) some test cases witness a mismatch between outputs of the C and the Rust code that the translator has created so far (line 10). We ensure this by providing a prompt to the discriminator to continue retrying until the generated test cases meet the above two requirements. Once successful, the newly generated test cases ($Batch$) are added to the full test set. The translator agent then takes the updated test cases as guidance to improve the translation. It is set up with prompts that repeatedly ask it to pass all the test cases in the current test set (lines 16-19). If the translator runs out of retries but still cannot fully pass the test set $T$, we let it skip and move to the next iteration. After a fixed maximum number of iterations, ACToR stops the loop and outputs the final translation and the test set generated.
As the iteration progresses, it becomes increasingly challenging for the discriminator agent to identify differences solely by examining the source code or conducting careful manual testing. To increase the limit, we design a simple agent-friendly fuzzing script to accelerate the mismatch discovery. With the help of a fuzzer script, the discriminator gains more power in finding errors in corner cases. We show the effectiveness of the fuzzing-augmented discriminator embodied in ACToR in evaluations.
4 Evaluation
The goal of a good defense is to retain the functionality of the given C code, and make the defended program fully memory-safe. Specifically, we evaluate ACToR on the following criteria:
  1. 1.
    Scalability. What size of C programs can ACToR automatically translate?
  2. 2.
    Correctness. Do the Rust translations maintain equivalent behavior to the C program? Do the Rust translations use safe Rust?
  3. 3.
    Ablation. How much does the adversarial agentic design contribute to correctness?
4.1 Experiment Settings
Table 1: Benchmark Statistics
(a) Micro Benchmark
Program LOC Funcs Cov.
printf 371 11 95%
expr 452 17 92%
fmt 416 12 91%
test 528 17 78%
join 475 10 90%
csplit 303 7 88%
Total 2545 74 89%
(b) Macro Benchmark by Category
Category LOC Funcs Count
Text Processing 13269 355 21
File Dir. Mgmt. 7465 151 15
Scripting Utils 4204 118 8
Sys. Info Status 2908 70 10
Env. Proc. Ctrl. 204 8 3
Total 28050 702 57
Benchmarks. We use two sets of benchmarks in our evaluation: a micro-benchmark set and a macro-benchmark set. The micro-benchmark comes from a prior user study on C to Rust translation, which contains 6 standalone C programs, with an average of 424 lines of code each. We manually craft at least 60 test cases for each program in the micro benchmark with an average of 89% line coverage, which are used to evaluate the correctness of translations. We use this small-scale benchmark to demonstrate ACToR’s ability to adapt to different LLMs and coding agent frameworks, and to perform an ablation study.
Our main macro-benchmark is a set of 57222The full set of BSDCoreUtils contains 69 programs, but we filter out 12 of them from our main results due to three reasons: (1) some utilities are trivial (e.g., true, yes) (2) certain utilities require special environments (e.g., stty) or privileges (e.g., sudo) not available in our experimental sandbox; and (3) certain utilities perform destructive operations (e.g., rmdir) that may break the testing harness. real-world system utility programs (e.g., ls, cp, xargs) obtained from a software project named BSDCoreUtils. It has an average of 492 lines of code each. We use this benchmark to evaluate the scalability of ACToR, and compare correctness to our baselines. These programs do not come with tests or test runners. We preprocessed these programs to add test runners, which can automatically run a set of given tests for each program, compare C and Rust outputs, as well as produce coverage reports. Note that while we provide the test runner, it is infeasible to manually create comprehensive test cases for all programs to evaluate which translation strategy is more correct. Thus, unlike micro-benchmarks, we perform a relative evaluation between ACToR and baselines, as described below in the Evaluation Metrics section.
To minimize measurement noise and make our experiments automatable, our evaluation focuses on single-threaded C programs that have deterministic behaviors across different executions, which is currently an assumption of our testing framework that both the translator and the discriminator depend on. Table 1(b) summarizes the statistics of two benchmarks, where we categorize the 57 programs into five categories: process execution control, file/directory management, shell scripting commands, system information access, and textural content processing.
Agent Frameworks and Models. We use Claude Code (Anthropic, 2025a) with Claude Sonnet 4 (Anthropic, 2025b) as the main agent framework and LLM for our experiments. To test if our approach generalizes to other agent frameworks and LLMs, we also performed experiments using Mini-SWE-Agent (SWE-agent, 2025), a popular open-source agent framework, with both GPT 5 mini (OpenAI, 2025) and Claude 4 Sonnet as the LLM. We use ‘ClaudeCode’, ‘SWE+Sonnet’, and ‘SWE+GPT5mini’ to represent them in later sections, respectively.
Comparison Baselines. We compare ACToR with two non-adversarial baselines, and one adversarial LLM agent baseline. (1) Naive baseline: a single agent instructed to translate the C program using only an initial set of seed tests, and does not iterate to add more tests. (2) Coverage baseline: a two-agent setup similar to ACToR, but the discriminator agent is only instructed to find additional tests that increase the line coverage of the C program, so the new tests do not specifically try to reveal inequivalent behavior. (3) ACToR-NoFuzz: ACToR but without the fuzzing tool. Comparing ACToR with the first two baselines demonstrates the benefit (if any) of our adversarial design. Comparing ACToR with ACToR-NoFuzz demonstrates the benefit (if any) of our proposed fuzzing tool.
Evaluation Metrics. Our primary evaluation metric to evaluate the correctness of an individual translation is pass rate, which is the percentage of tests that pass for a given set of tests. For our micro-benchmark, we use our manually written test cases to compute pass rates for each benchmark. For our macro-benchmarks, which do not have manually written tests, we compute pass rates using a set of tests produced by a competing method, and refer to this metric as relative pass rate. For example, when comparing ACToR with the coverage baseline, we compute pass rate for ACToR using the tests generated by the coverage baseline, and vice versa. We use this approach because a given method might get biased towards its own tests (usually a given method passes all of its own tests). Relative pass rate provides a strong signal for comparing ACToR against our baselines. In addition, the test sets used to compute relative pass rate achieve an average of 89.7% line coverage on the micro benchmark and 90.6% line coverage on the macro benchmark, by unioning test cases across translation methods. This indicates they are a useful signal for overall correctness as well.
Experimental Configuration. ACToR and all baselines start with 15 manually crafted, diverse seed tests based on the C code for each program  [^3] The seed tests carry intended test format that both translator and discriminator of ACToR use. This avoids a cold-start for ACToR as well as all other systems we compare to.. For the coverage baseline, ACToR-NoFuzz, and ACToR, we run 10 iterations of test generation and translation. On each iteration, we generate 3 new tests, resulting in a total of 45 tests per program in the end. We choose 10 iterations because more iterations usually do not result in improvements in pass rate and limit the incurred financial costs. We generate 3 tests per iteration because larger values lead to a decline in quality of tests, and smaller values have slower convergence. In considering cost, we allow translator agents and discriminator agents to proceed directly to the next iteration if they are unable to make progress (i.e., generate valid test cases or correct the translation to pass the current tests) after 3 retries. The prompt for the translator agent is shared among all the translation methods and put in Appendix B. The prompts for the test generator agent in the coverage baseline and the discriminator of ACToR are in Appendix B.
4.2 Evaluation Result on Micro Benchmark
Figure 2: Overall correctness (pass rate) achieved by ACToR on micro benchmark across different settings compared with naive baseline, on 3 different agent-model settings.
Correctness & Adaptability. We first test ACToR on the micro benchmark with different agents and LLMs. To analyze the correctness improvement, we compare ACToR approach with a naive baseline on pass rate. The experiment results per program are shown in Figure 2. All translations from ACToR and the naive baseline are 100% safe and free of unsafe sections. ACToR consistently improves the translation correctness on all 3 agent-model settings. With Claude Code, the validation pass rate rises from an average of 79.3% (one-shot naive baseline) to 92.1% after 10 iterations. For Mini-SWE-Agent, pass rates increase from 81.9% to 90.7% with Sonnet-4 and from 84.1% to 86.8% with GPT-5 mini on average. Across agent-LLM settings, the closed-source Claude Code achieves the highest final pass rate. We therefore use Claude Code + Claude-Sonnet-4 as the default configuration for macro benchmarks.
Ablation Study. We conduct an ablation to know how the adversarial design in ACToR drives correctness gains on the micro-benchmark. We analyze the improvement offered by two design choices: (a) having an adversarial discriminator in the workflow to guide the translator, and (b) forcing the discriminator agent to use the fuzzing script we provided. We compare three variants of ACToR: coverage-baseline, ACToR-NoFuzz, and the full ACToR (with fuzzing). The results summed across programs at the final iteration are reported in Figure 3. Entry $(\text{row},\text{column})$ shows the relative pass rate of the row method on the column method’s tests. All three methods largely pass their own tests as expected.
Figure 3: The relative comparison among 3 different translation methods on Claude Code. Entry $(\text{row},\text{column})$ is the relative pass rate of the row’s translation on the column method’s tests.
On coverage-baseline tests, ACToR-NoFuzz and the full ACToR reach 92% and 88% passing rates, respectively; in the reverse direction, coverage-baseline achieves less than 70% passing rate on tests from either ACToR variant. This indicates that adversarial discriminators produce tests that more effectively expose semantic mismatches and result in more correct translations. Between ACToR variants, the full ACToR fails more on its own tests, but attains 82% vs. 75% when cross-tested against ACToR-NoFuzz. This suggests that the discriminating agent uses the fuzzing script to surface more challenging bugs in the translated code. We also measure the final test coverage of each method on the C code. The result is shown at the bottom of Figure 3. The coverage baseline works as expected and achieves the highest line coverage on the C code than two other ACToR variants. However, it is worse in correctness. This further supports the conclusion that adversarial translation is more efficient in improving correctness compared with purely sampling new test cases based on coverage.
On our micro benchmark, the full ACToR system substantially improves translation correctness compared with the naive baseline. The adversarial discriminator is crucial to improving correctness, and the fuzzing tool further strengthens the discovery of mismatches.
Figure 4: The relative pass rate when cross-comparing ACToR and coverage baseline (Cov-Base) on Claude Code at iteration 10. For each program, the left bar shows evaluating the translation from ACToR on tests generated in coverage-baseline; the right bar is evaluating the translation from coverage-baseline on ACToR tests. The length of each program in LoC is presented next to the program name.
4.3 Evaluation on Macro Benchmark
We compare ACToR results with the coverage baseline to examine the improvement achieved in correctness, using the relative pass rate metric. We use the Claude Code agent and Claude-Sonnet-4 for this experiment. In considering the necessary system calls used in the translated code, we discourage the use of unsafe by instructing in prompts, but do not enforce full safe Rust in the middle of the translation loop. We enforce 100% safe Rust at the end of the translation in ACToR, as an additional simple post-processing agent call.
The cross-comparison result at the $10^{th}$ iteration [^4] On 3 programs ‘sync‘, ‘users‘, and ‘mknod‘, both translation methods abort before 10 iterations after multiple reruns. We therefore compare them on the maximum available iteration. is shown in Figure 4. The left bar for each program is the relative pass rate of testing ACToR’s translation on the test cases from the coverage baseline. The right bar shows the reverse: relative pass rate of the coverage baseline against ACToR’s test cases. We see that ACToR outperforms the coverage baseline on 55/57 programs on the relative pass rate. ACToR gets 100% relative pass rate on 26 programs when running on the tests from the coverage baseline, while the coverage baseline can only get full pass on 1 program. On 57 programs, ACToR achieves an average 93.9% relative pass rate, 18.9% higher than the 75.0% relative pass rate obtained by the coverage baseline. The cross-comparison confirms that adversarial agentic design in ACToR plays a significant role in improving the translation correctness. In an absolute (non-relative) sense of correctness, by taking the union of the test cases from both ACToR and the coverage baseline, we get a test set of an average 90.6% coverage. ACToR achieves a 95.3% pass rate on it. This demonstrates the utility of ACToR on translating real-world projects with high correctness. We further analyze the only benchmark program, specifically the pr program, for which ACToR has lower correctness than the coverage baseline in the Appendix A.
ACToR automatically translates the BSDCoreUtil benchmark with high correctness, outperforming the coverage baseline on 5555/$57$ programs and achieving an average of $93.9\%$ relative pass rate.
5 Conclusion
We present ACToR, an agent-based automatic translation approach that aims for functionally correct C to Rust translations. Our method introduces a collaborative pipeline in which two agents iteratively refine the translations. At each iteration, the discriminant agents detect the semantic mismatches between the source code and the translated program, and report them as failing test cases. The translator agents then improve the translation to pass the current test set. Our experiments demonstrate the power of ACToR’s adversarial agentic setup to produce translations with high correctness, surpassing the quality of translation as compared to non-adversarial setups. A noteworthy advantage that ACToR brings to the translation landscape is the potential for large-scale translation with minimal or no human intervention, making ACToR the first approach of this kind.
Reproducibility Statement
The artifact containing detailed instructions, benchmark programs, and evaluation scripts will be released soon. Section 3 describes the algorithmic design of ACToR, including assumptions and iteration parameters, while Section 4 outlines our evaluation methodology and benchmark datasets. Appendix B includes the prompt template used for the translator and discriminator agents. Together, these resources enable independent verification of both our experimental setup and results.
Appendix A Case Study
ACToR outperformed coverage baseline on 55/57 programs in the macro benchmark, while failing on one program name ‘pr‘. We further analyze the reason behind this failure. The correctness gap is caused by the corner difference between the environment where we run ACToR and the validation environment when we test the translation, due to an imperfect testing harness. The translator agent claims that all tests were passed during the translation procedure, but actually got timeouts on most cases during validation. After manually adjusting the code (1 line edit) to adapt to the validation environment, the translation achieves an $84\%$ relative pass rate, higher than the coverage baseline.
(ACToR/ Coverage Baseline) Translator Agent Prompt Task Description: ‘‘‘ Please translate the C program in the current folder to safe Rust (no unsafe code). Please create a xxx_rust folder in the current folder, where xxx is the name of the C program (you can see from the Makefile). Please use Cargo to build the Rust program. Once you have the compiled Rust program (typically in target/debug/xxx‘), please run testcmp.sh to compare the output with the original C program. Please run testcmp.sh --help to understand how to use the test script. Please debug to make sure the translation passes all the tests. ‘‘‘ Please start the translation now.
Figure 5: The task prompt for the translator agent.
(Coverage Baseline) Test Generator Agent Prompt — Adding Tests Task Description: ‘‘‘ Please add more tests for the C program in the current folder. The goal is to increase the line coverage while keeping tests valid per our definition. GOALS OF ADDING TESTS: 1. Increase the line coverage. 2. Avoid duplication with existing tests; focus on new, meaningful scenarios and edge cases. 3. Ensure validity: Run ‘./testcmp.sh compare ./xxx and confirm the C program passes all tests. Internally, this means two compiled versions of the C program (‘./xxx.ref and ‘./xxx‘) must produce identical results, ensuring issues are fixable (not nondeterminism). RESOURCES FOR ADDING TESTS: - C source code files: ‘*.c and ‘*.h for understanding intended behavior. - Makefile for building (‘make all‘) and cleaning (‘make clean‘). - testsXX.jsonl‘: existing tests and format examples. - testcmp.sh for running comparisons: ‘./testcmp.sh compare <BINARY>‘. To get the coverage of tests, run ‘./testcmp.sh coverage‘. - $x$.gcov file after collecting coverage, to check which lines are not covered. TASK: Add 3 new test cases to testsXX.jsonl where $X$ is the largest index among all testsXX.jsonl files. <...Test formatting requirements> <...Detailed documentation for the input format of the test runner> ‘‘‘ Now please add 3 more tests as requested above.
Figure 6: The task prompt for the test generator agent of the coversge baseline.
(ACToR) Discriminator Agent Prompt — Adding Tests Task Description: ‘‘‘ Please add more tests for the C program in the current folder. The goal is to reveal behavioral differences between the current Rust translation and the C reference while keeping tests valid per our definition. GOALS OF ADDING TESTS: 1. Expose divergences between Rust and C behavior (Rust should fail where C passes). 2. Avoid duplication with existing tests; focus on new, meaningful scenarios and edge cases. 3. Ensure validity: Run ‘./testcmp.sh compare ./xxx and confirm the C program passes all tests. Internally, this means two compiled versions of the C program (‘./xxx.ref and ‘./xxx‘) must produce identical results, ensuring issues are fixable (not nondeterminism). RESOURCES FOR ADDING TESTS: - C source code files: ‘*.c and ‘*.h for understanding intended behavior. - Makefile for building (‘make all‘) and cleaning (‘make clean‘). - testsXX.jsonl‘: existing tests and format examples. - testcmp.sh for running comparisons: ‘./testcmp.sh compare <BINARY>‘. To get the coverage of tests, run ‘./testcmp.sh coverage‘. IF {\$allow_fuzz} - fuzzer_template.py for fuzzing: python fuzzer_template.py <BINARY1> <BINARY2>‘. You can read its code to understand how to generate tests. ENDIF - Rust translation project if present (e.g., ‘<prog>_rust with Cargo). Build with cargo build and try ‘./testcmp.sh compare <path_to_rust_binary>‘ to probe differences. - $x$.gcov file after collecting coverage, to check which lines are not covered. TASK: Find 3 new input cases that expose mismatches between C code and Rust translation. Add these 3 new test cases to testsXX.jsonl where $X$ is the largest index among all testsXX.jsonl files. <...Test formatting requirements> IF {\$allow_fuzz} <...Details on the usage of the fuzzer script> ENDIF <...Detailed documentation for the input format of the test runner> ‘‘‘ Now please add 3 more tests as requested above.
Figure 7: The task prompt for the discriminator agent of ACToR. The usage of the fuzzing tool is controlled by ‘\$allow_fuzz‘ in the prompt.
Appendix B Prompt Template
For Mini-SWE-Agent, we keep the system prompt that describes the agent workflow unchanged as the original one during all the experiments. For Claude Code, we begin directly with the task described in the user message, without any additional text preceding it. The task prompt structure for the translator agent is shown in Figure 5. This prompt is shared among all translation methods during evaluation. The prompt structure for the test generation agent of the coverage baseline is shown in Figure 6. The prompt for the discriminator agent for ACToR is shown in Figure 7. The fuzzing mode is controlled by the ‘allow_fuzz‘ section in the prompt.